Calculate the Medicare levy payable (if any) for the year ending 30 June 2020 of the following amounts of taxable income for a single Australian resident taxpayer:

(a) $15,000 (b) $23,000 (c) $40,000 (d) $140,000
Answer #1

a.

Not required to pay any Medicare levy since the txable income is less than the threshold for medicare levy purposes

Lower threshold limit for medicare levy is $22,801.

b.

Taxable income falls between the lower and upper threshold for medicare levy purpose.

Lower threshold limit for medicare levy is $22,801.

upper threshold limit for medicare levy is $28,801.

Medicare levy payable = 10% of the of the taxable income in excess of lower threshold = ($23,000-$22,801)* 10% = $199 * 10% =  $19.90

c.

Taxable income is higher than the upper threshold.

So, Medicare levy = 2% of the of the taxable income = $40,000 * 2% = $800

d.

Taxable income is higher than the upper threshold.

Medicare levy = 2% of the of the taxable income = $140000 * 2% = $2,800

[For this case, medicare surcharge also applicable since taxable income is more than the threshold limit of $90,000 for MLS purpose. So, surcharge = ( 105000-90000 ) * 1% + (140000-105001) * 1.25% = $150+ $437.50 = $587.50 ]